[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Nov 17, 2015

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Tue Nov 17 12:47:04 UTC 2015


DX AWIPS2 Hardware Refresh on Tuesday (TODAY). It will not effect
operations.



HSU:

The hurricane hot line appears to be working normally.

HSU reported the it took longer than normal to produce some Atlantic ATCF
graphics. Monica or Craig will investigate.




TAFB:

Chris fixed the SREF model ingest procedure in GFE.

12Z SFC analysis failed to reach OPC. Andy experienced this several times
over the past week. Re-sending using trop.launch fixed the problem each
time. TSB will investigate.

Other Issues:

 The VM server (nhc-lx-kvm01) that hosts NHC's backup websites will be
rebooted Thursday to resolve a memory issues.




-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20151117/2050315f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list